# WWAggr: A Window Wasserstein-based Aggregation for Ensemble Change Point Detection

### v1

@@ V-E3 Choice of calibration method
Table VI: Expected Calibration Error (see [41], we want to minimize $\downarrow$ it) values for the BCE models trained on the Explosions dataset before and after calibration. The best metric value is highlighted in bold.
@@ V-E4 Choice of probabilistic distance
In this experiment, we examine the sensitivity of the proposed aggregation technique from Algorithm 1 to the choice of the probabilistic distance $d(\cdot)$. Although all three considered options ($\hat{W}_{1}$, $\hat{W}_{2}$, MMD) produce close detection quality when the alarm threshold is carefully selected, the 1-Wasserstein distance behaves much more stable with 1 — 5 standard predefined thresholds (see Figure 4). Thus, this is the most empirically appropriate distance variant for WWAggr.
@@ VI Limitations and future work
The main limitation of our approach is its computational complexity. On the Explosions dataset, the average GPU inference time per sequence for BCE/InDiD increases from $40.5\pm 1.2$ ms for a single model to $400.0\pm 3.2$ ms for an ensemble of size $K=10$ with “naive” aggregations. Our WWAggr method with the same ensemble size took  $417.1\pm 5.0$ ms, incurring a minor computational overhead ($\approx 5\%$) compared to naive approaches. WWAggr has the potential to overcome this difference and be even more efficient if the sliced [47] or projected [48] Wasserstein distances are used.
This work targets high-dimensional data, where assumption-free deep CP detectors are the only reasonable solution [6, 13]. However, for lower dimensions, the increase in computational cost for ensembles of such deep detectors may outweigh its modest performance benefits. In this case, one could resort to Bayesian Neural Networks [49], although their efficacy strongly depends on making proper prior choices [17]. Another alternative for low-dimensional data could be to ensemble classical statistical methods, such as CUSUM [24]. While we hypothesize our aggregation strategy would also be effective for these procedures, classical methods eliminate the calibration of CP scores and, consequently, any guarantee for easy alarm threshold choice. Thus, we leave these ideas for future investigation.
@@ VII Conclusion
In our experiments, ensembles of deep CPD detectors consistently outperform standalone models, even with simple aggregation methods like averaging. However, these “naive” procedures are sub-optimal for the change point detection setting. Thus, we introduce WWAggr, the novel ensemble prediction aggregation procedure based on the Wasserstein distance. This procedure significantly boosts detection quality by capturing the peculiarities of the CPD task, being model-agnostic and suitable for various supervised and unsupervised CPD base models. As a result, WWAggr advances an ensemble, improving the detection $F_{1}$-score on the video data by up to $20\%$ compared to basic aggregations.
To ensure robustness against alarm threshold selection, we apply model calibration, making the proposed approach more practical. Consequently, in inference, almost any reasonable threshold value results in a decent performance.
@@ Appendix A Dataset descriptions
Here, we provide a detailed description of the datasets used in experiments. The dataset summary statistics are presented in Table III.
@@
As the most complicated datasets, we use high-dimensional semi-structured video data. To construct these samples, real-world RGB videos were taken from the UCF-Crime anomaly detection dataset [46], and their CPD markup was provided by [6]. In the Explosions dataset, CPs correspond to any explosions and fire shot by a CCTV camera. ”Normal” videos do not include any emergencies. These situations follow the CPD problem statement as the explosion’s consequences (fire, smoke, etc.) are seen for some time afterwards. Clips from Road Accidents follow a similar pattern but are more challenging. In this case, CPs indicate car crashes on the streets. As traffic accidents do not necessarily split the video into two regimes, we use this dataset to measure the range of applicability of CPD models. In these two video datasets, every clip has a length of $16$ frames, and every frame is a tensor of size $240\times 320\times 3$.
@@ Appendix B Implementation details
We conducted all the experiments from the paper using Python 3 and the PyTorch framework. The code is available online [^2] https://anonymous.4open.science/r/wwagr-ensemble-cpd-E7FF. All the models were trained on two NVIDIA GeForce GTX 1080 Ti GPUs. For the implementation of base deep CPD models, which form various ensembles, we mostly followed the original works and inherited the proposed model architectures. Below, we describe the models’ architecture choice and the training pipeline hyperparameters. Additionally, we provide details on implementing the proposed sliding-window Wasserstein-based aggregation procedure.
@@ B-A Details on supervised ensembles
We utilize two sequence-to-sequence supervised approaches. Both models employ recurrent LSTM-based [50] architectures but are trained with different loss functions. The first one, denoted as BCE, is a generic neural network that considers CPD as a classification task and directly optimizes the standard binary cross-entropy loss [1, 6, 34]. In contrast, the second InDiD model leverages the principles of the CPD task and optimizes the detection delay and number of false alarms [6]. Both approaches are recognized for their high quality across various real-world CPD tasks, including the detection of changes in video samples. We refer the reader to the original paper for more details.
These BCE and InDiD models consist of LSTM [50] blocks followed by dense layers. To stabilize this model, we added Layer Norm to the LSTM cells, which resulted in more robust training and better final metrics compared to the original paper [6]. The rest of the hyper-parameters and the training pipeline were taken from the repository [^3] https://github.com/romanenkova95/InDiD.
@@ B-B Details on unsupervised ensembles
Note that both types of unsupervised base learners (TS-CP2 and SN-TS2Vec) are based on representation learning. Consequently, they detect change points by computing the cosine distances between subseries representations, and these distances, bounded to $[-1,1]$, serve as CP scores $p_{t}$. On the other hand, supervised baselines operate with CP probabilities $p_{t}\in[0,1]$. In order to align these two approaches, we suggest post-processing TS-CP2 and SN-TS2Vec outputs with the following transformation: $\tilde{p}_{t}=\left(1-p_{t}\right)\cdot\mathbb{I}\left(p_{t}\geq 0\right)$, where $\mathbb{I}(\cdot)$ represents the indicator function. As a result, the final CP scores lay in the range from 0 to 1, allowing us to apply WWAggr in all the cases.
B-C Change point detection for video data

### v2

@@ V-E3 Choice of calibration method
Table VI: Expected Calibration Error (see [42]; should be minimized $\downarrow$) values for the BCE models trained on the Explosions dataset before and after calibration. The best metric value is highlighted in bold.
@@ V-E4 Choice of probabilistic distance
In this experiment, we examine the sensitivity of the proposed aggregation technique from Algorithm 1 to the choice of the probabilistic distance $d(\cdot)$. Although all three considered options ($\hat{W}_{1}$, $\hat{W}_{2}$, MMD) produce close detection quality when the alarm threshold is carefully selected, the 1-Wasserstein distance behaves in a more stable way with 1 — 5 standard predefined thresholds (see Figure 4). Thus, this is the most empirically appropriate distance variant for WWAggr.
@@ VI Limitations and future work
The main limitation of our approach is its computational complexity. On the Explosions dataset, the average GPU inference time per sequence for BCE/InDiD increases from $40.5\pm 1.2$ ms for a single model to $400.0\pm 3.2$ ms for an ensemble of size $K=10$ with “naive” aggregations. Our WWAggr method with the same ensemble size took  $417.1\pm 5.0$ ms, incurring a minor computational overhead ($\approx 5\%$). WWAggr has the potential to overcome this difference and be more efficient if the sliced [48] or projected [49] Wasserstein distances are used.
Additionally, this work targets high-dimensional data, where assumption-free deep CP detectors are the only reasonable solution [6, 13]. However, for lower dimensions, the increase in computational cost for ensembles of such deep detectors may outweigh its modest performance benefits. In this case, one could resort to Bayesian Neural Networks [50], although their efficacy strongly depends on making proper prior choices [17]. Another alternative could be to ensemble classical statistical methods, such as CUSUM [25]. While we hypothesize our aggregation strategy would also be effective for these procedures, classical methods eliminate the calibration of CP scores and, consequently, any guarantee for easy alarm threshold choice. Thus, we leave these ideas for future investigation.
@@ VII Conclusion
In our experiments, ensembles of deep CPD detectors consistently outperform single models, even with simple aggregation methods like averaging. However, these “naive” procedures are sub-optimal for the change point detection setting. Thus, we introduce WWAggr, the novel ensemble prediction aggregation procedure based on the Wasserstein distance. This procedure significantly boosts detection quality by capturing the peculiarities of the CPD task, being model-agnostic and suitable for various supervised and unsupervised CPD base learners. As a result, WWAggr advances an ensemble, improving the detection $F_{1}$-score on the video data by up to $20\%$ compared to basic aggregations.
To ensure robustness against alarm threshold selection, we apply model calibration, making the proposed approach more practical. Consequently, in inference, almost any reasonable threshold value results in a decent performance quality.
Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement with Skoltech №139-10-2025-033.
@@ Appendix A Dataset descriptions
As more complicated datasets, we use high-dimensional semi-structured video data. To construct these samples, real-world RGB videos were taken from the UCF-Crime anomaly detection dataset [47], and their CPD markup was provided by [6]. In the Explosions dataset, CPs correspond to any explosions and fire shot by a CCTV camera. ”Normal” videos do not include any emergencies. These situations follow the CPD problem statement as the explosion consequences (fire, smoke, etc.) are seen for some time afterwards. The clips from Road Accidents follow a similar pattern but are more challenging. In this case, CPs indicate car crashes on the streets. As traffic accidents do not necessarily split the video into two regimes, we use this dataset to measure the range of applicability of CPD models. In these two video datasets, every clip has a length of $16$ frames, and every frame is a tensor of size $240\times 320\times 3$.
@@ Appendix B Implementation details
We conducted all the experiments from the paper using Python 3 and the PyTorch framework. The code is available online [^1] https://github.com/stalex2902/wwagr-ensemble-cpd. All the models were trained on two NVIDIA GeForce GTX 1080 Ti GPUs. For the implementation of base deep CPD models, which form various ensembles, we mostly followed the original works and inherited the proposed model architectures, which is describes below. Additionally, we provide details on implementing the proposed sliding-window Wasserstein-based aggregation procedure.
@@ B-A Details on supervised ensembles
We utilize two sequence-to-sequence supervised approaches. Both models employ recurrent LSTM-based [51] architectures but are trained with different loss functions. The first one, denoted as BCE, is a generic neural network that considers CPD as a classification task and directly optimizes the standard binary cross-entropy loss [1, 6, 35]. In contrast, the second one, InDiD model, leverages the principles of the CPD task and optimizes the detection delay and number of false alarms [6]. Both approaches are recognized for their high quality across various real-world CPD tasks, including the detection of changes in video samples. We refer the reader to the original paper for more details.
These BCE and InDiD models consist of LSTM [51] blocks followed by dense layers. To stabilize this model, we added Layer Norm [52] to the LSTM cells, which resulted in more robust training and better final metrics compared to the original paper [6]. The rest of the hyper-parameters and the training pipeline were taken from the repository [^2] https://github.com/romanenkova95/InDiD.
@@ B-B Details on unsupervised ensembles
Note that both types of unsupervised base learners (TS-CP2 and SN-TS2Vec) are based on representation learning. Consequently, they detect change points by computing the cosine distances between subseries representations, and these distances, bounded to [1,1][-1,1], serve as CP scores $p_{t}$. On the other hand, supervised baselines operate with CP probabilities $p_{t}\in[0,1]$. In order to align these two approaches, we suggest post-processing TS-CP2 and SN-TS2Vec outputs with the following transformation: $\tilde{p}_{t}=\left(1-p_{t}\right)\cdot\mathbb{I}\left(p_{t}\geq 0\right)$, where $\mathbb{I}(\cdot)$ represents the indicator function. As a result, the final CP scores lay in the range from 0 to 1, allowing us to apply WWAggr in all the cases.
B-C Video data preprocessing